OpenStudy (anonymous):

The numerator is 1 more than the denominator. If the numerator and the denominator are both increased by 2, the fraction will be 1 over 4 less than the original fraction. Find the original fraction.

OpenStudy (whpalmer4):

original fraction could be \[\frac{x+1}{x}\]the new fraction is then \[\frac{x+1+2}{x+2} = \frac{x+3}{x+2}\]and we know that \[\frac{x+1}{x} - \frac{1}{4} = \frac{x+3}{x+2}\]only some tedious algebra separates you from the answer :-)